### v1

# Digital Network Twins for Next-generation Wireless: Creation, Optimization, and Challenges

Abstract
Digital network twins (DNTs), by representing a physical network using a virtual model, offer significant benefits such as streamlined network development, enhanced productivity, and cost reduction for next-generation (nextG) communication infrastructure. Existing works mainly describe the deployment of DNT technologies in various service sections. The full life cycle of DNTs for telecommunication has not yet been comprehensively studied, particularly in the aspects of fine-grained creation, real-time adaptation, resource-efficient deployment, and security protection. This article presents an in-depth overview of DNTs, exploring their concrete integration into networks and communication, covering the fundamental designs, the emergent applications, and critical challenges in multiple dimensions. We also include two detailed case studies to illustrate how DNTs can be applied in real-world scenarios such as wireless traffic forecasting and edge caching. Additionally, a forward-looking vision of the research opportunities in tackling the challenges of DNTs is provided, aiming to fully maximize the benefits of DNTs in nextG networks.
I Introduction
Although digital twin (DT) has been successfully applied in the fields of manufacturing and complex system operation and maintenance, research related to the creation, deployment, and application of DT for next-generation (nextG) networks is still in its infancy. In this regard, digital network twins (DNTs) emerge as a promising solution [1], offering a multitude of significant benefits, such as accelerated network development processes, enhanced resource utilization, easy management, and faster networking innovation with reduced costs. Specifically, in the network deployment phase, the availability of DNT can deliver a nextG network planning and operation tool to see the entire life cycle of network and management details before physically placing wireless sites. If system enhancement is desired, performance parameters can be adjusted and simulated in a DNT without imperiling the operation of the entire infrastructure. Then, during the network operation phase, DNT can enhance the access and analysis of already deployed network systems to enable remote commissioning and diagnostics, which lowers service costs by remotely configuring faulty parts of a network facility, e.g., ordered and replaced accordingly for the network operator or other involved researchers; Lastly, in the network design phase, a DNT can also be used to predict network performance and behavior in complex environments, design newer and better systems from the learned history of infrastructure operating conditions, and optimize facilities’ efficiency and output.
However, before its full realization, representing a physical network using a virtual model faces many challenges: 1) how to build a high-fidelity twin involving various physical objects, communication entities, signal propagation phenomenon, and network attributes synergistically; 2) how to adaptively evolve the built DNTs taking into account the highly dynamic nature of nextG networks; 3) how to integrate the DNT model with the existing network infrastructure and systems while considering communication and computational constraints; and 4) how to gather the representative data from a DNT and its physical counterparts for assisting various downstream applications.
To resolve these practical challenges, several surveys and tutorials about DNTs have been proposed [2, 3, 4, 5, 6, 7, 8, 9], mainly describing the deployment of DNT technologies in various service sectors. [10, 3, 4] present a broad survey to explore the integration of DTs into networks for potential application scenarios, such as aviation, healthcare, 6G networks, and intelligent transportation systems. [5] focuses on deploying DNTs at edge to improve performance such as user throughput and security, and reducing the cost of communication, computation, and caching. [6, 8] outline the recent trends in machine learning and explore their integration into DNTs for efficient network management. [7] analyzes the current state of the DNT paradigm and classifies the potential threats associated with it, considering its functionality layers and the operational requirements to ensure the trustworthy use of a DT. Specifically, [9] takes more complex information into account, capturing and representing network dynamics as well as stimulating with the modeled reality in real-time. Nevertheless, despite the widespread interest in deploying DNTs for telecommunication, it has not yet been comprehensively studied about the full lifecycle of DNTs, particularly in fine-grained creation, real-time adaptation, resource deployment optimization, and security.
This article presents a broad overview of DNTs and unveils the full potential of their integration into communication networks, covering the fundamentals, the emergent applications, and critical challenges in multiple dimensions. We start by exploring challenges and opportunities in DNT model creation and adaption, in the aspects of mapping solutions, twinning data generation, and real-time synchronization requirements. Then, we consider the joint optimization of mapping and resource management to deploy the DNTs in practical network systems, by integrating split learning, graph neural networks, and proximal policy optimization within a reinforcement learning framework. Moreover, the security issues within the DNTs are unveiled, with a particular focus on adversarial attacks and information leakage. Lastly, we showcase concrete application scenarios of DNTs supported by use-case experiments. The paper concludes by outlining future research directions in addressing the above technical challenges.
Figure 1: The differences between traditional DT and digital network twins (DNTs). DNTs go beyond by coordinating and managing the transitions between multiple DTs within a network to optimize overall performance in real time.
II Preliminaries and Overview
II-A Digital Twin vs. Digital Network Twin
A digital twin (DT) is a virtual representation of a physical system designed to reflect its runtime process and critical properties. It covers the entire lifecycle of each object, continuously updating from real-time data, and employs reasoning to predict future dynamics and inform decision-making. As an extension, DNTs offer advanced capabilities beyond traditional DTs by incorporating network-specific elements. While DTs enable continuous data exchange between isolated digital models and physical systems, DNTs extend this by jointly modeling, simulating, and optimizing complex communication networks in real time. DNTs simulate network behaviors such as traffic patterns, latency, and data transmission efficiency while adapting to dynamic network topologies and ensuring quality of service (QoS). They address multi-dimensional challenges such as inter- and intra-connectivities, security, and resource management more effectively than DTs, which typically focus on individual systems. A key feature of DNTs is their ability to manage transitions between multiple twins within a network and dynamically allocate resources—such as bandwidth and computing power—across interconnected twins. This is critical in large-scale, dynamic networks where traffic, connectivity, and user demands fluctuate. DNTs ensure efficient data flow, resilience, and adaptability across the network, coordinating interactions between multiple twins to optimize performance.
As shown in Fig. 1, a physical network (PN) consists of user devices (i.e., virtual reality headsets, sensors, servers, base stations (BSs), etc), network protocols, and communication channels. Specifically, a user device is defined as a physical network object (PNO) that can be described by a set of internal properties (i.e., physical dimension, location, mobility patterns) and a set of external properties (i.e., data transmission, access channel). In particular, external properties are used to capture how PNO affects other PNOs in the PN. On the other hand, the DNT, consisting a network of twins, is a real-time virtual mapping of the PN such that it is expected to comprise the same components with the same internal and external properties. Specifically, the DNT is built based on the following two processes: 1) Mapping all PNOs with their internal and external properties as well as their interrelations into a digital environment; 2) Self-evolving this digital environment according to the changes in any internal or external properties of PNOs in the PN. For example, a user device switches its data transmission from BS1 to BS2 in the PN, which will also affect the external properties (e.g., spectrum access) of other user devices due to the varying communication interference. Therefore, when the DNT updates the external property for a mobile device, it must also update the properties for other affected user devices accordingly. This considered representation model will guide the fundamental designs of DNT creation, updates, and applications as discussed in the following sections. Notably, while the existing works [10, 5, 6, 7, 8, 9] extensively explored the concept and applications of DNT architectures, such as simulating 5G networks and optimizing spectrum usage, practical challenges regarding efficient creation and real-time updates remain underdeveloped. For instance, a conceptual survey on DNTs in [10] highlights how DTs enhance network performance through real-time monitoring and predictive analytics, emphasizing their role in intelligent network management. Another recent survey [7] explores the security landscape of DTs by categorizing potential threats and vulnerabilities in various downstream applications. It mainly emphasizes the increasing attack surface introduced by DTs, particularly due to their integration with physical and cyber domains. Orthogonal to existing literature, this article specifically addresses the foundational building blocks and online adaptation of DNTs within communication networks, introducing novel mapping mechanisms and resource deployment strategies to address a network of DTs. We also investigate continuous monitoring and optimization to mitigate security and privacy concerns in critical network settings.
Figure 2: Summary of the challenges and opportunities of the DNT.
II-B DNT Applications
II-B1 What-if Analysis and Troubleshooting
DNT offers a unique safe sandbox to simulate various runtime network statuses and configurations, where the impact and potential loss of dangerous operations can be safely evaluated and analyzed. In this context, DNT not only facilitates the training of network operators in countering emergent service disruptions and cybersecurity issues but also serves as a virtual environment for the detection of network misconfiguration, performance bottleneck, and link failures without any impact on the physical network. Notably, proactive measures powered by DNT excel conventional reactive measures for network fault recovery, which often suffers from a longer delay to prevent lasting severe economic loss. For example, the prediction of poor link quality for a mobile device availing URLLC service, at a certain point in the future (time or distance) will enable the network to prepare for such an event with proactive measures such as allocating more resources, switching to different frequency bands, preparing for new beams, etc.
II-B2 Network Planning and Management
The increasing scale and complexity of modern networks pose challenges in meeting the demands of larger networks efficiently. By utilizing DNT, ISPs can predict when existing networks will exhaust their resources given a specified growth in users. Furthermore, these performance estimates offer invaluable insights for planning optimal upgrades capable of handling such growth effectively. In essence, the DNT empowers network operators to make informed decisions and anticipate the need for network upgrades. Notably, a DT prototype developed within the Hexa-X initiative [11] introduces dynamic control of 5G slicing, ensuring that QoS guarantees are upheld for each slice, catering to diverse traffic patterns. It also facilitates resource allocation for each slice in accordance with service level agreements (SLAs). Furthermore, the DNT envisions equitable capacity assignment and real-time planning for future 6G infrastructure [8], underscoring its adaptability and foresight in shaping the landscape of upcoming DT-assisted network technologies.
III Challenges and Opportunities
This section addresses the challenges, mitigation strategies, and prospects of DNTs, as summarized in Fig. 2. We begin by examining the creation and adaptation of DNTs, with a particular emphasis on optimizing operational efficiency. Next, we explore their practical application in wireless resource management, where the continuous monitoring, assessment, and stabilization of DNT quality at scale pose significant challenges. Lastly, we highlight the security risks that emerge when integrating machine learning models into DNTs, especially in privacy-sensitive and mission-critical cyber infrastructures, acknowledging these as critical concerns.
III-A Creation and Adaptation of Digital Network Twin
III-A1 Vertical Mapping from Physical Networks
One major evolution of nextG mobile and wireless networks is being driven by a move to higher carrier frequencies, e.g., millimeter-wave (mmWaves) and Terahertz (THz). Different from the lower-frequency radio, the diffraction ability of an mmWave/THz signal is much weaker and less reliable due to its smaller wavelength, which means that deployed objects in the environment have a significant impact on communication attributes and performance. This yields insight into deriving network attributes from environment details, including objects’ locations, sizes, material types, and reflective properties. Thus, instead of cloning all details from the physical network to DNT, which is inefficient and even impractical, one possible solution is to map essential environment information from the physical networks, and then derive the corresponding attributes in the DNT locally through an environment-aware mapping approach as in [12]. This necessitates employing machine learning frameworks like long short-term memory (LSTM) and convolutional neural networks (CNNs) to effectively capture spatial and temporal domain knowledge within a wireless network. Synchronous federal learning (FL) is particularly effective for this vertical mapping because it requires all participating devices to update the global model simultaneously. This simultaneous updating process ensures high consistency across devices and improves overall model convergence, leading to more accurate and stable outcomes at this initial stage. Specifically, multiple twins can be created using a periodic clustering mechanism that virtualizes network regions covered by different base stations, taking into account their distinct geographical and wireless characteristics.
III-A2 Horizontal Mapping for DNT update
Different from the vertical mapping approach that gathers the information from physical environments to build a network twin, one primary goal of horizontal mapping is to reuse and migrate the existing DNTs and then create a new one when necessary (as shown in Fig. 3 with the blue lines). To this end, two fundamental approaches can be exploited, including: a) GNN-based rendering: graph neural network (GNN) can be used to evolve existing DNT for environment dynamics. This includes detecting the variations of the physical network and then updating the DNT accordingly. Given the output of the GNN, an unsupervised network representation learning model [13] can be trained to determine whether to preserve the current DNT attributes and structure and then update the corresponding nodes according to the joint vertex embeddings when necessary. b) Multi-twin merging: When mapping the physical network into the digital space, we may need to create multiple DNTs that fulfill different tasks via transfer learning and/or game theory, e.g., a radio-map twin used for network planning and resource allocation, a channel-knowledge twin for CSI acquisition and signal processing, etc. This involves the process of selecting appropriate twins among a set of available candidates, and then fusing the corresponding attributes in spatial and/or temporal domain. Another objective is to continuously update the existing twin model. Asynchronous federated learning (FL) is particularly suited for this phase, as it permits devices to contribute updates to the global model at different times. This flexibility enhances scalability by allowing devices to participate despite varying availability, while also supporting different computational capacities, making the process more adaptable to diverse environments.
III-A3 Twinning Data Generation
One main challenge of utilizing AI/ML methods to construct DNTs is the necessity to generate a substantial volume of representative data capturing both environmental intricacies and coupled network attributes. However, this task is often impractical and unfeasible in real-world networks due to the extensive measurement campaigns required. Thus, one potential solution is to synthetically generate high-quality data covering a wide range of fine-grained network scenarios, which is then used to develop a faithful, efficient approach for the DNT creation. For instance, a quasi-deterministic (Q-D) ray tracer [12] can be used to generate various instance data within a known scenario covering different access point (AP)/user locations and protocol settings. Such a model can capture the geometrical properties of the channel for each transceiver and generate the profile of delay, path gain, angle of departure (AoD), angle of arrival (AOA), etc, for every path component. However, the data quality from the ray tracer is highly dependent on the abstracted channel model, which needs to be well-designed to acquire an accurate channel impulse response. In this regard, a small amount of measurement data from real-world experimental testbeds will contribute to the data generation process in two aspects: a) help derive the aligned scenario distributions, and b) create a feedback path to tune the related deterministic channel parameters such that the ray tracer can produce a more accurate dataset that matches the actual measurements.
Figure 3: ML-assisted DNT mapping framework.
III-B Wireless Resource Management for Digital Network Twins
Different from current resource management work that optimizes temporal and spatial allocations only for the physical network, resource management herein must consider networking resource (i.e., spectrum, transmit power and computational power) management and mapping method selection to optimize the performance of both physical networks and DNTs. In particular, each mapping method requires unique physical network data to generate DNTs, thereby resulting in different data transmission requirements. For example, to map communication attributes between two entities, horizontal mappings require the digital counterpart to learn the network evolution while vertical mappings mainly require environment parameters from the physical counterpart. In this regard, joint optimization of mapping method selection and resource management yields the following unique challenges.
III-B1 DNT Quality Assessment
To optimize the wireless resource management of DNT and the physical network, the first challenge is to design a novel performance metric that can capture how the mapping strategy, data collection, and transmission affect DNT creation and optimization over the actual network. Different from the current DT work that only maps physical objects into virtual space, DNT requires mapping network environment details and underlying network attributes into twins. Here, one promising method is to leverage multi-attribute utility theory – a powerful multiple-criteria decision analysis tool that can effectively analyze multiple dependent and correlated variables in a joint utility function. Different from current multi-attribute utility functions that use weight parameters to summarize different metrics, which may not be applied for dependent metrics, multi-attribute utility theory builds a joint utility function that can capture the effects of all network attributes, protocols, and physical objects along the DNT creation, analogous to the generation of a joint probability distribution in probability theory, thereby enabling the analysis of the correlations among dependent network attributes and parameters.
III-B2 Optimized DNT Deployment
Another challenge is to address the optimization of the DNT performance (e.g., efficiency and accuracy of generation process) via determining mapping methods (i.e., horizontal or vertical mappings introduced in Subsection III A) for each network entity and wireless resource (i.e., bandwidth, transmit power, beamforming parameters) for data transmission from physical networks to DNTs. Since mapping method selection and actual wireless resource allocation are always coupling and time-dependent, where the current mapping and resource allocation will explicitly affect future decisions, one may study the use of reinforcement learning (RL) to maximize the joint utility function derived from the quality assessment stage.
However, implementing RL for joint mapping selection and wireless resource management encounters two main challenges. First, representing the dynamics of twining objects and network attributes as a state in an RL framework is non-trivial, as nextG networks may involve numerous object nodes with diverse potential dynamics, which cannot be adequately captured by a limited-size vector. One promising solution is to use a GNN framework with heterogeneous graphs to represent the network dynamics due to its ability to represent different physical networks and DNTs with a unique embedding vector and to analyze the inherent relationships among them. The second challenge lies in how to deal with the continuous action space over the actual network (e.g., with bandwidth and power allocation, beamforming process). This is because standard RL (e.g., Q-learning) that uses value functions to select optimal actions cannot deal with continuous action space. While discretizing these optimization variables will sacrifice the DNT performance, this motivates the use of proximal policy optimization (PPO) that can stabilize and output an optimal action according to an RL state.
III-B3 Large-scale Twin Optimization
The training and implementation of a centralized RL framework for DNT optimization require the collection of all physical network dynamics and DNT states thus probably leading to high communication overhead and implementation cost, particularly in a large-scale nextG wireless network. As a result, the RL model must be extended to a distributed scheme, referred to as DRL, which enables distributed APs and user devices in a physical network to cooperatively determine their mapping methods and resource utilization thus efficiently optimizing the utility function. To this end, one solution is to explore a split learning (SL) approach to decide on how to divide GNNs and RL models across different network entities, and how to leverage that division to share data and perform collaborative RL for the creation of DNTs. Unlike standard SL, which categorizes neural networks based solely on data distribution and learning performance, DNT uses separated neural network models that take into account factors such as computational resources, transmit power, wireless spectrum resources, and the dynamics of each device. Given the divided neural networks, the next step is to investigate different information-sharing procedures among DNTs and analyze their impacts on collaborative RL performance.
III-C Security of Inter-Realm Communication with DNTs
While numerous state-of-the-art studies are centered on crafting and implementing DT systems across diverse domains, there remains a notable dearth of emphasis on safeguarding and enhancing the security and privacy aspects of these immersive, mixed realms. Specifically, the nascent DNTs and their interfaces to the physical network infrastructure might become potential targets for malicious activities. This underlines the emergence of less-explored security and privacy risks, among which are adversarial attacks that engender counterfeit representations, alongside the vulnerability of information leakage across distributed wireless devices and apps. The intricate interplay and added complexity of the DNT ecosystem give rise to less-recognized but impactful threats, especially in the following areas.
III-C1 Adversarial attacks during mapping and synchronization
First, unauthorized data manipulation during the digital mapping process undermines trust in virtual services by empowering attackers to exploit manipulated digital models for deceptive services. Second, continuous synchronization can open doors to inter-realm adversarial attacks, wherein subtle data perturbations may result in a persistent divergence between the digital and physical systems. Effectively countering data poisoning attacks throughout the digital mapping and synchronization process while preserving digital representation accuracy remains a challenging and under-explored task. To tackle this problem, potential solutions involve mitigating security risks posed by malicious actors attempting to manipulate the digital model at an early stage and establishing a new security framework against the inter-realm adversarial attack, such as continuously paring the critical digital network components with the physical counterparts to identify whether the mapping data are trustworthy or not, followed by a mask-predict based countermeasure to drop out or replace the poisoned data/model pieces while maintaining the freshness of digital representation.
III-C2 Information leakage with ML models in DNTs
Given that ML models form a core component of the DNTs, a significant threat model pertains to ML parameter leaks within the network. This involves potential adversaries like VR users and servers excluded from trustworthy training processes. These adversaries, whether active or passive, pose the risk of inferring raw data from both VR users and servers. The consequential risk is the potential leakage of sensitive information, encompassing user details as well as system-level operational details when leveraging these ML models in virtual environments. One potential solution is to design a novel distributed learning framework that enables distributed wireless devices to collaboratively train ML models while preventing the leakage of sensitive data of connected users during the training of ML models for DNT generation and synchronization. This framework combines quantization theory and federated learning, to enable wireless devices to efficiently generate a globally optimal ML model and prevent information leakage without sacrificing ML performance.
III-C3 Privacy risks within DNTs
The creation and use of DNTs raise significant privacy concerns due to the vast amount of data they capture, process, and store. These concerns encompass unauthorized access, misuse of sensitive information, surveillance, and data breaches. DNTs, being intricate and interconnected, risk inadvertently disclosing personal internet usage, preferences, and intellectual property to unauthorized parties. To address these concerns, robust data encryption and access control measures can be implemented to ensure secure and authorized access to the digital environment. Privacy-by-design principles can also mitigate risks by integrating privacy protections into the DNT mapping phase, including data minimization strategies during physical data collection, which balances the twin fidelity and system security. Additionally, employing techniques like decentralized federated learning among multiple split twins [14] can enhance privacy by distributing model updates instead of raw data to a central physical network controller, and/or with differential privacy adding noise to updates to prevent reverse engineering from DNTs to the physical networks.
IV Case Studies
This section showcases the creation and optimization of DNTs in two case studies: wireless traffic forecasting and reliable edge caching, with respect to the proposed methods in Sec. II-A and II-B.
IV-A Joint Vertical-Horizontal Twinning for Traffic Prediction
In this section, we introduce the use of DNT for wireless traffic forecasting, where BSs can be clustered and virtually mapped into DNTs based on their distinct geological and network characteristics. As explained in Sec. III-A, these DNTs can be constructed by vertical mapping (V-twin) and horizontal mapping (H-twin) for DNT initialization and evolution, respectively. Here, V-twin exploits a synchronous federated learning framework to initialize a global twin model from the historical traffic data of virtualized network clusters. H-twin, on the other hand, is implemented with an asynchronous federated learning scheme that dynamically updates distributed twin models in response to network or environmental changes. The resulting self-evolved twins can then be applied for collaboratively forecasting precise time-series data, such as heterogeneous wireless traffic in cellular networks.
Figure 4: Mapping accuracy and deployment costs.
Specifically, we employ the real-world network traffic dataset [15] of Milan City, which is composed of urban telecommunications behaviors intricately divided into 10,000 grid cells, with each cell equipped with a dedicated BS that captures a wealth of time-series data records, including detailed information on short message service (SMS), call activities, and internet usage patterns. We adopt mean squared error (MSE), mean absolute error (MAE), and normalized root mean square error (NRMSE) as key performance metrics to measure the distinctions between ground-truth and generated data from the mapped DNTs. The ground truth data are time-series real-world traffic data collected from the physical cellular network. Another key performance metric is the cost of the mapping process, which includes both training and communication time. In traditional DNT, all mapping is performed on a centralized server, requiring raw data to be uploaded for one-time modeling, which significantly increases deployment costs compared to our joint V-H twin method with continuous adjustments. H-twin incurs higher costs compared to V-twin, as DNTs demand continuous updates over extended periods rather than just initial construction. For comparison, we skip the BS clustering stage to create two twinning baselines, namely, V-twin w/o. cluster. and H-twin w/o. cluster. As shown in Fig. 4 with an example of SMS traffic data, all four twinning methods provide comparable prediction error rates, with the H-twin achieving slightly lower prediction errors. However, H-twin needs to aggregate local twin models from the BSs and dispatch them back, thereby slightly increasing the communication cost. Notably, clustering of BSs can significantly reduce the sum of communication and training costs by approximately 43.8% for H-twin. Despite the overhead introduced by H-twin, its asynchronous twinning process can effectively handle real-time heterogeneous wireless data, resulting in a reliable and accurate DNT as the time-series data evolves.
IV-B DNT-Assisted Reliable Edge Caching
DNTs can enhance edge caching optimization by embedding intervention modules into data-driven models for greater stability. Integrating safe RL algorithms with constraint modules, DNTs can be applied to optimize and safeguard caching decisions within the RL framework in a three-fold manner: 1) simulate network behavior and predict caching outcomes through accurate vertical mapping with historical data; 2) act as reliability nets, supporting risk-free exploration and policy testing in controlled environments (e.g., avoid overloading some BSs); and 3) generate rare events, allowing RL agents to train on sparsely distributed corner cases by producing synthetic yet realistic data across diverse conditions (e.g., maintain the hit rate in scenarios with highly dynamic content demand).
Figure 5: DNT facilitates reliable edge caching by accounting for reliability constraints and minimizing risks.
Figure 6: DNT with reliable RL balances loads among BSs.
Here, we consider a cellular wireless network consisting of five BSs. Each BS is equipped with a local cache unit with a capability of 150 cache slots. A BS can only access its local cache unit and is forbidden to retrieve any cache units from its neighboring BSs. Each BS provides service for eight clients. The service can be overlapped, e.g., one client can be served by up to two BSs. We then create a DNT via vertical mapping from historical caching data, including information about user requests, contents, frequency, etc. The DNT serves as a data generator, providing ample datasets for training and predictive evaluation of different cache replacement policies. It produces one-step forecasts for the subsequently requested content. Moreover, various sets of historical caching data can be employed to feed the DNT, generating datasets that cover both common and rare wireless scenarios. We then implement four edge caching systems with or without the DNT. 1) RL, a basic reinforcement learning model that uses data collected from physical cellular networks for training and prediction solely; 2) RL w/ DNT, the model trained with data from both BSs and the virtual DNT; 3) Reliable RL, the model augmented by integrating safety state, action and reward modules to monitor and reduce the load imbalance among BSs; 4) Reliable RL w/ DNT, the reliable RL model trained with diverse data from both BSs and the virtual DNT. Figs. 5 and 6 evaluate the performance of four systems in terms of caching hit rate, number of risk interventions (introduced by safe modules from reliable RL), and BS load. The inclusion of DNT with safety modules achieves a higher hit rate with balanced BS loads, hence enhancing the reliability and efficacy in managing cached content across densely deployed small-cell BSs. The use of DNT also reduces the intervention from the elaborated safety constraints, suggesting that the augmented data from DNTs can improve the robustness of RL models with fewer interventions.
V Conclusion
This article presents a broad overview of DNTs and unveils the vast potential of their integration for nextG wireless networks. We start by introducing the basic concepts and fundamental design of DNTs, with three highlighted applications including what-if analysis and network planning. We then describe the critical research challenges and opportunities of DNTs, in the aspects of creation, adaptation, optimization, and security. Lastly, we showcase the use of DNTs in two wireless scenarios: traffic forecasting and edge caching. We present a variety of research opportunities in addressing the challenges of DNTs, with the goal of fully harnessing their potential in future communication networks.